# Unveiling Entity-Level Unlearning for Large Language Models: A Comprehensive Analysis

### v2

@@ 4.2 Effect of Unlearning Steps
Additionally, we observe that during unlearning, all methods show a more pronounced decline in performance on the other entities set compared to the other non-target sets including the rerain set, world fact set and real authors set, which are introduced by pre-training. This phenomenon suggests that entity knowledge introduced through fine-tuning may be more susceptible to unlearning. Based on this observation, we hypothesize that this vulnerability arises because the information introduced through fine-tuning primarily affects the model’s behavior rather than enhancing its knowledge. To test this hypothesis, we further compare the performance of the other entities set to that of the other non-target sets in both pre-training and fine-tuning scenarios.
@@
We present the unlearning performance of the five algorithms on pre-trained and fine-tuned entities, as illustrated in Figure 5. For the pre-trained entities, all five methods display a consistent trend between the other entities set and the other non-target sets. In contrast, for the fine-tuned entities, the Grad. Diff., KL Min., and Pref. Opt. algorithms cause more significant damage to the other entities set than the other non-target sets. These findings suggest that the knowledge introduced during fine-tuning is more vulnerable to unlearning interventions, confirming our hypothesis about its fragility compared to the original pre-trained knowledge. Based on this analysis, future research should explore more robust methods for injecting target knowledge in entity-level unlearning tasks, aiming to better simulate the pre-trained knowledge which constitutes a significant majority of the knowledge within LLMs.
@@ 5 Related Work
5.1 Methods of LLM Unlearning
LLM unlearning (Liu et al., 2024b) has attracted rising attention owing to its potential to address privacy (Jang et al., 2022) and security (Barrett et al., 2023) concerns. Current unlearning algorithms for LLMs (Cao and Yang, 2015; Yu et al., 2023; Jang et al., 2022) focus on minimizing the impact of undesirable training data while preserving the integrity of other model knowledge based on the Forget Set and Retain Set. Specifically, these methods (Jang et al., 2022; Chen and Yang, 2023) apply Gradient Ascent (GA) on the Forget Set to unlearn and add additional auxiliary loss on the Retain Set akin to gradient descent (Liu et al., 2022) and KL minimization (Maini et al., 2024) to mitigate undesired effects. Additionally, inspired by the alignment capabilities of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022), several researchers adopt preference optimization methods to refine model outputs, like Direct Policy Optimization (DPO) (Rafailov et al., 2024) and Distributional Dispreference Optimization (D2O) (Duan et al., 2024). Moreover, some researchers utilize model editing methods to remove sensitive information by adjusting knowledge-related parameters (Wu et al., 2023; Wang et al., 2024b). However, it has been proven that the deleted content can be reverse-engineered from the edited model (Patil et al., 2023).
@@ 5.2 Evaluations of LLM Unlearning
Recent research has introduced several benchmarks and tasks for LLM unlearning from various aspects (Ji et al., 2024; Li et al., 2024; Maini et al., 2024; Lynch et al., 2024; Jang et al., 2022). Among these, the Weapons of Mass Destruction Proxy (WMDP) benchmark (Li et al., 2024) specifically targets dangerous knowledge in biosecurity, cybersecurity, and chemical security. While these task settings primarily address the forgetting of explicit instances, they pay less attention to entity-level unlearning, which involves completely forgetting an entity. Notably, Eldan and Russinovich (2024) explored a particular task to unlearn the entity "Harry Potter", but it is proven that the knowledge is not entirely erased from the unlearned model (Shi et al., 2023). Maini et al. (2024) presented TOFU consisting of 200 fictitious author profiles to assess unlearning methods from Model Utility and Forget Quality. However, this benchmark only focused on ideal scenarios with exact forget sets. Similarly, the RWKU benchmark (Jin et al., 2024) chooses 200 real-world famous people as entities for unlearning, with a more practical task setting. Nonetheless, it still focuses on specific pieces of knowledge related to target entities rather than the entire entity, making it closer to the instance-level unlearning task. To sum up, there is a notable absence of systematic analysis and evaluation for entity-level unlearning.
@@ 6 Conclusion
In this paper, we propose a novel task: entity-level unlearning for LLMs, which is required in many practical scenarios. We evaluate trending unlearning algorithms in this work and reveal that existing unlearning methods struggle to effectively erase the entire entity from the target model. Furthermore, we find that the size and knowledge coverage of forget set play crucial roles in the performance of the algorithms. Additionally, our analysis shows that entities introduced through fine-tuning are more vulnerable to unlearning compared to pre-trained entities, underscoring the need for more robust entity injection techniques. These findings offer valuable insights and encourage future research to develop corresponding unlearning algorithms and explore more precise knowledge-probing methods to achieve improved entity removal.
@@
Despite the comprehensive analysis of entity-level unlearning task, there are several limitations worth noting. First, our work primarily focuses on analyzing the entity-level unlearning task and identifying potential insights for improvement, rather than proposing a specific unlearning algorithm, which will be a focus of our future research. Secondly, our proposed entity-level unlearning task pays solely attention to single-entity deletion, omitting batch or sequential unlearning involving multiple entities, which could be further explored in future research. Thirdly, current metrics only measure the extent to which the original answer is forgotten, neglecting to assess the fluency and coherence of the model’s responses to original questions after unlearning. Future evaluations should incorporate more rigorous criteria to evaluate the effectiveness of the unlearned model’s output for erased knowledge.
@@ Appendix C Details for Analysis
C.1 Analytical Experiments on Probing Set
When analyzing the impact of probing sets on the unlearning effect, we introduce a novel metric, Knowledge Coverage, designed to assess the overlap in knowledge between a probing set and a target set. Specifically, for each QA pair in the probing set, we compute the BERTScore against every QA pair in the target set, retaining the highest score and its corresponding target set index. It is possible for QA pairs in the target set to achieve multiple BERTScores, meaning they may be the best match for several QA pairs in the probing set. Knowledge Coverage is then calculated as the ratio of the sum of the maximum BERTScores for each sample in the target set to the total number of QA pairs in the target set. The $KC_{score}$ can be written as:
@@ C.1 Analytical Experiments on Probing Set
Where $B(\cdot)$ is the function to calculate the BERTScores between the probing set $S_{P}$ and the target set $S_{T}$.
For all five algorithms, we conduct analytical experiments on the probing set size, using AdamW with warm-up during the first epoch, a batch size of 2, a learning rate of $10^{-5}$, and 6 epochs.
Additionally, for all five algorithms, we also conduct analytical experiments on the Knowledge Coverage of the probing set, using AdamW with warm-up during the first epoch, a batch size of 4, a learning rate of $10^{-5}$, and 12 epochs. Specially, we create multiple probing sets by substituting different ratios of QA pairs from the target set, while keeping the total size constant at 20, to examine the impact on Knowledge Coverage. The Knowledge Coverage of each probing set is shown in Table 7.
@@
Table 7: Knowledge Coverage of the probing sets obtained by replacing different percentages of QA pairs from the target set.

### v3

@@ 4.2 Effect of Unlearning Steps
Additionally, we observe that during unlearning, all algorithms show a more pronounced decline in performance on the other entities set compared to the other non-target sets including the retain set, world fact set and real authors set, which are introduced by pre-training. This phenomenon suggests that entity knowledge introduced through fine-tuning may be more susceptible during unlearning. Based on this observation, we further compare the performance of the other entities set to that of the other non-target sets in both pre-training and fine-tuning scenarios.
@@
We present the unlearning performance of the five algorithms on pre-trained and fine-tuned entities, as illustrated in Figure 5. For the pre-trained entities, all five methods display a consistent trend between the other entities set and the other non-target sets. In contrast, for the fine-tuned entities, the Grad. Diff., KL Min., and Pref. Opt. algorithms cause more significant damage to the other entities set than the other non-target sets. These findings suggest that the knowledge introduced during fine-tuning is more vulnerable to unlearning interventions, confirming our hypothesis about its fragility compared to the original pre-trained knowledge. Based on this analysis, future research should explore more robust methods for injecting target knowledge in entity-level unlearning tasks, aiming to simulate better the pre-trained knowledge which constitutes a significant majority of the knowledge within LLMs.
@@ 5 Related Work
5.1 Algorithms of LLM Unlearning
Current unlearning algorithms for LLMs (Cao and Yang, 2015; Yu et al., 2023; Jang et al., 2022) focus on minimizing the impact of undesirable training data while preserving the integrity of other model knowledge based on the Forget Set and Retain Set. Specifically, these methods (Jang et al., 2022; Chen and Yang, 2023) apply Gradient Ascent (GA) on the Forget Set to unlearn and add additional auxiliary loss on the Retain Set akin to gradient descent (Liu et al., 2022) and KL minimization (Maini et al., 2024) to mitigate undesired effects. Additionally, inspired by the alignment capabilities of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022), several researchers adopt preference optimization methods to refine model outputs, like Direct Policy Optimization (DPO) (Rafailov et al., 2024) and Distributional Dispreference Optimization (D2O) (Duan et al., 2024). Moreover, some researchers utilize model editing methods to remove sensitive information by adjusting knowledge-related parameters (Wu et al., 2023). However, it has been proven that the deleted content can be reverse-engineered from the edited model (Patil et al., 2023).
@@ 5.2 Evaluations of LLM Unlearning
Recent research has introduced several benchmarks and tasks for LLM unlearning from various aspects (Ji et al., 2024; Li et al., 2024; Maini et al., 2024; Lynch et al., 2024; Jang et al., 2022). Among these, the Weapons of Mass Destruction Proxy (WMDP) benchmark (Li et al., 2024) specifically targets dangerous knowledge in biosecurity, cybersecurity, and chemical security. While these task settings primarily address the forgetting of explicit instances, they pay less attention to entity-level unlearning, which involves completely forgetting an entity. Notably, Eldan and Russinovich (2024) explored a particular task to unlearn the entity "Harry Potter", but it is proven that the knowledge is not entirely erased from the unlearned model (Shi et al., 2023). Maini et al. (2024) presented TOFU consisting of 200 fictitious author profiles to assess unlearning methods from Model Utility and Forget Quality. However, this benchmark only focused on ideal scenarios with exact forget sets. Similarly, the RWKU benchmark (Jin et al., 2024) chooses 200 real-world famous people as entities for unlearning, with a more practical task setting. Nonetheless, it still focuses on specific pieces of knowledge related to target entities rather than the entire entity, making it closer to instance-level unlearning.
@@ 6 Conclusion
In this paper, we propose a novel task, entity-level unlearning for LLMs, which is required in many practical scenarios. We evaluate popular unlearning algorithms on this task and reveal that existing unlearning methods struggle to effectively erase the entire entity from the target model. Furthermore, we find that the knowledge coverage and size of forget set play crucial roles in the performance of the algorithms. Additionally, our analysis shows that entities introduced through fine-tuning are more vulnerable than pre-trained entities during unlearning, underscoring the need for more robust entity injection techniques. These findings suggest promising directions for future research, encouraging the development of more precise knowledge-probing methods and corresponding unlearning algorithms to enhance entity removal.
@@
Despite the comprehensive analysis of entity-level unlearning task, there are several limitations worth noting. First, our work primarily focuses on analyzing the entity-level unlearning task and identifying potential insights for improvement, rather than proposing a specific unlearning algorithm, which will be a focus of our future research. Secondly, the entity-level unlearning task in our work pays solely attention to single-entity deletion, omitting batch or sequential unlearning involving multiple entities, which could be further explored in future research. Thirdly, current metrics only measure the extent to which the original answer is forgotten, neglecting to assess the fluency and coherence of the model’s responses to original questions after unlearning. Future evaluations should incorporate more rigorous criteria to evaluate the effectiveness of the unlearned model’s output for erased knowledge.
@@ Appendix C Details for Analysis
C.1 Analytical Experiments on Forget Set
When analyzing the impact of forget sets on the unlearning effect, we introduce a novel metric, Knowledge Coverage, designed to assess the overlap in knowledge between a forget set and a target set. Specifically, for each QA pair in the forget set, we compute the BERTScore against every QA pair in the target set, retaining the highest score and its corresponding target set index. It is possible for QA pairs in the target set to achieve multiple BERTScores, meaning they may be the best match for several QA pairs in the forget set. Knowledge Coverage is then calculated as the ratio of the sum of the maximum BERTScores for each sample in the target set to the total number of QA pairs in the target set. The $KC_{score}$ can be written as:
@@ C.1 Analytical Experiments on Forget Set
Where $B(\cdot)$ is the function to calculate the BERTScores between the forget set $S_{F}$ and the target set $S_{T}$.
For all five algorithms, we conduct analytical experiments on the size of the constructed forget set, using AdamW with warm-up during the first epoch, a batch size of 2, a learning rate of $10^{-5}$, and 6 epochs.
Additionally, for all five algorithms, we also conduct analytical experiments on the Knowledge Coverage of the constructed forget set, using AdamW with warm-up during the first epoch, a batch size of 4, a learning rate of $10^{-5}$, and 12 epochs. Specially, we create multiple constructed forget sets by substituting different ratios of QA pairs from the target set, while keeping the total size constant at 20, to examine the impact on Knowledge Coverage. The Knowledge Coverage of each constructed forget set is shown in Table 7.
@@
Table 7: Knowledge Coverage of the constructed forget sets obtained by replacing different percentages of QA pairs from the target set.